Skip to main content

IODP Expedition 376

Brothers Arc Flux


Daily Science Report for 19 May 2018

Location: Hole U1527C (34°51.6625′S, 179°3.2534′E, water depth 1464 mbsl; proposed Site NWC-1A)

Science Update: Half-length Cores 16R to 20R penetrated from 214.0 to 238.0 m and recovered 8.5 m (35%). We pumped mud sweeps for hole cleaning at various depths. After cutting Core 20R at 0700 h, we observed a tight hole and had to work it from 234 to 125 m with high drill string torque and poor hole conditions, leading to stuck pipe at 125 m (30 m below the end of the casing string). We pumped mud sweeps at 234, 145, and 130 m. At 2045 h, we offset the vessel ~75 m to access a drill pipe connection at the rig floor to retrieve Core 20R and release the bit in an attempt to free the drill string. We offset the vessel to the original position at 2230 h and continued to work the stuck drill string from 125 to 116 m with high torque.

Cores 14R to 20R have been split and described today. They consist of pervasively altered, poorly sorted mono- to polymict lapilli-tuff to tuff-breccia, being either matrix- or clast-supported. Maximum cobble-sized, subrounded to subangular volcanic clasts show clay and chlorite alteration. Vugs are filled with zeolite, silica, and sulfide. The latter also occurs disseminated. Braided massive magnetite-pyrite and silica veins are rarely present. Cores 14R to 20R continue the highly to completely hydrothermally altered pyroclastic deposit recovered in Cores 11R to 13R.


Daily Science Report for 18 May 2018

Location: Hole U1527C (34°51.6625′S, 179°3.2534′E, water depth 1464 mbsl; proposed Site NWC-1A)

Science Update: Cores 2R to 15R penetrated from 99.9 to 214.0 m. The first nine of these cores partially took only 10 min to cut and had no to extremely poor recovery (0 to 0.54 m), retrieving pebbles of fresh to slightly altered, sparsely to highly plagioclase-pyroxene phyric, dark gray dacite lava with moderate to high vesicularity. While cutting Core 11R, we encountered a substantial formation change at ~187 m and core recovery increased to 29%. The time to cut Core 11R increased to 70 min. To enhance the core recovery even more, we decided to cut half-length cores (4.8 m), and recovery for the remainder of today’s cores improved to 72%. We circulated 410 barrels of mud sweeps for hole cleaning at various depths. The change in formation is reflected in the lithology—consolidated matrix-supported volcaniclastic breccia. Cores 11R to 13R have been split and described, revealing at least three different types of lapilli- to block/bomb-sized volcanic clasts (altered plagioclase phyric dacite, altered medium-grained clasts, fresh fine-grained clasts). The breccia is pervasively altered, displaying a yellowish brown to greenish blue color with secondary clay, iron oxide, zeolite, and subordinate pyrite.


Daily Science Report for 17 May 2018

Location: Hole U1527C (34°51.6625′S, 179°3.2534′E, water depth 1464 mbsl; proposed Site NWC-1A)

Science Update: Today we successfully drilled the 94 m long casing string into the seafloor when the pilot bit had reached 99.9 m. Immediately after this was completed, we deployed the go-devil to activate the hydraulic release tool, which freed the casing at 0045 h. We then started pulling the drill string out of the hole, reaching the seafloor at 0315 h. We set back the top drive, retrieved the subsea camera system, and the bit returned to the rig floor at 1150 h. After taking apart the drilling assembly, we flushed the mud motor and underreamer with fresh water. We then started making up the rotary core barrel (RCB) bottom-hole assembly at 1300 h and lowered it to the seafloor. We deployed the subsea camera system and reentered Hole U1527C at 2014 h. We picked up the top drive, dropped a core barrel, and recovered the subsea camera system. At midnight, the end of the casing string was reached and we began RCB coring from 99.9 m. Scientists continued to analyze and prepare reports on their Hole U1527A results. A scientist gave a talk on the alteration mineralogy of magmatic-hydrothermal systems studied at the surface of Brothers volcano to date.


Daily Science Report for 16 May 2018

Location: Hole U1527C (34°51.6625′S, 179°3.2534′E; water depth 1464 mbsl, proposed Site NWC-1A)

Science Update: We continued to recondition the drilling assembly and reentry system for our second attempt to drill it into the seafloor. We picked up the 94 m long casing string, reassembled the reentry funnel in the moonpool, welded the funnel and hard rock landing frame to the casing, inserted the drilling assembly, and lowered it to the seafloor. We deployed the subsea camera system at 0945 h to observe the reentry system while drilling in the casing, and we started drilling Hole U1527C at 1130 h. At midnight, the bit had reached 96.7 m with only a few more meters to achieve total depth. Scientists continued to collect, analyze, and write up Hole U1527A data. They also performed Kuster Flow-Through Fluid Sampler contamination tests in preparation for planned borehole fluid sampling.


Daily Science Report for 15 May 2018

Location:
• Hole U1527B (34°51.6519′S, 179°3.2526′E, water depth 1464 mbsl; proposed Site NWC-1A)
• Hole U1527C (34°51.65′S, 179°3.25′E, water depth 1464 mbsl; proposed Site NWC-1A)

Science Update: We continued drilling the 94 m long casing string into the seafloor until the pilot bit reached the final planned depth of 105.5 m at 0030 h. We activated the hydraulic release tool (HRT) to detach the drilling assembly from the casing at 0114 h, and we began recovering the drill string and the subsea camera system. When we attempted to pull the HRT through the moonpool, we observed that the reentry system was still attached to the drill string. We landed and dismantled the reentry system in the moonpool, and we were able to free the drilling assembly underreamer arms from the casing. When the drilling assembly was pulled up to the rig floor at 1530 h, we discovered that the pilot bit was missing. We started to recondition the drilling assembly and reentry system so that we can install it in Hole U1527C, 20 m south of Hole U1527B. Scientists continued to collect, analyze, and write up Hole U1527A data.


Daily Science Report for 14 May 2018

Location: Hole U1527B (34°51.6519′S, 179°3.2526′E, water depth 1464 mbsl; proposed Site NWC-1A)

Science Update: We continued to drill in casing from 34.3 to 102.1 m. We had to work several tight spots from 82 to 102 m. Scientists continued to collect, analyze, and write up Hole U1527A data. Portable X-ray fluorescence measurements confirmed the dacitic composition of the volcanics recovered in Hole U1527A.


Daily Science Report for 13 May 2018

Location: Hole U1527B (34°51.64′S, 179°3.22′E, water depth 1464 mbsl; proposed Site NWC-1A)

Science Update: We spent the entire day continuing casing operations, consisting of (1) assembling a mud motor, underreamer, and 9⅞ inch tri-cone bit; (2) connecting the hydraulic release tool to the 10¾ inch casing assembly; and (3) attaching the reentry funnel and hard rock landing frame to the casing. At 1000 h, we started lowering the entire 94 m long casing string and drilling assembly to the seafloor. We deployed the subsea camera system at 1530 h to observe the reentry system while drilling in the casing. We began drilling the casing into the seafloor in Hole U1527B at 1715 h. At midnight, the end of the casing had reached 34.3 m. Scientists continued to collect, analyze, and write up Hole U1527A data.


Daily Science Report for 12 May 2018

Location:
• Hole U1527A (34°51.6528′S, 179°3.2397′E, water depth 1464 mbsl; proposed Site NWC-1A)
• Hole U1527B (34°51.64′S, 179°3.22′E, water depth 1464 mbsl; proposed Site NWC-1A)

Science Update: We continued RCB coring from 87.0 to 101.4 m with extremely poor recovery. We pulled the drill string out of the hole at 0635 h, set back the top drive, and the bit was back on the rig floor at 1120 h, ending Hole U1527A. Overall, only 1.27 m of core was recovered from the 101.4 m cored in Hole U1527A. We offset the ship 20 m to the east and started preparing the rig floor for assembling a reentry cone and 94 m of casing that we will drill into the seafloor. Scientists continued to collect, analyze, and write up their final Hole U1527A data. The volcanic rock recovered at Hole U1527A probably represents dacite. The Co-Chief Scientists gave a short overview of the seafloor rock types and mineralization studies in caldera wall proximal to Hole U1527B.


Daily Science Report for 11 May 2018

Location: Hole U1527A (34°51.6528′S, 179°3.2397′E, water depth 1464 mbsl; proposed Site NWC-1A)

Science Update: We continued RCB coring throughout the day. Cores 2R to 12R penetrated from 9.7 to 87.0 m and recovered 1.27 m. The drill string occasionally became stuck, which required frequent pumping of mud sweeps. Core recovery (1.6%) was extremely poor in the unconsolidated volcanic deposits. The material recovered is mainly fresh to slightly altered, sparsely plagioclase phyric volcanic rock of gray to dark gray color, showing moderate to high vesicularity. Subordinate, polymict pumiceous lapilli tephra was also recovered.


Daily Science Report for 10 May 2018

Location:
• In transit
• Hole U1527A (34°51.65′S, 179°3.21′E, water depth 1464 m; proposed Site NWC-1A)

Science Update: After a 246 nmi transit from Auckland, we arrived at Site U1527 at 0800 h. Following lowering thrusters, we put together the rotary core barrel (RCB) bottom-hole assembly, and started running it to the seafloor in preparation for coring. We attached an APCT-3 temperature shoe to the subsea camera system and conducted a survey of the seafloor; we located the precruise site survey seafloor marker at 1745 h. After completing the survey of six potential hole locations, we retrieved the camera system, picked up the top drive, and reassembled the rig floor. We began coring in Hole U1527A from a water depth of 1464 m at 2240 h, pumping tracer material and measuring downhole temperatures through readings of temperature strips contained in a housing tool attached to the RCB core barrel head. By midnight, Core U1527A-1R was retrieved with no recovery. After a Site U1527 kickoff meeting, the science party broke out into laboratory groups to continue preparing their laboratories and methods. An inspection of the third-party Petrospec spool-in thermocouple assembly was conducted, including a ship-to-shore videoconference meeting to familiarize the petrophysics group and IODP JRSO technical staff with the equipment. The science party started their shifts at 1200 h.


Daily Science Report for 9 May 2018

Location:
• Wynyard Wharf, Port of Auckland, New Zealand
• In transit to Site U1527 (proposed Site NWC-1A)

Science Update: The science party was given a tour to the helideck by the Schlumberger logging engineers, where they received an introduction to the downhole logging tools and their preparation for high-temperature downhole measurements at Brothers volcano. The Education and Outreach Officers provided the scientists an overview of their plans for the expedition. In the afternoon, the captain held the first fire and boat safety drill. The Expedition Project Manager (EPM) gave a talk on downhole measurements. The IODP JRSO Operations Superintendent and CDEX technical engineers introduced drilling and coring operations. The Publications Specialist introduced the science party to IODP publications, as well as their obligations during the expedition and afterwards. The second half of the science party was given a core flow tour by the EPM. The laboratory groups continued learning their laboratories and preparing instruments and methods. Upon completion of fueling in the early morning, the pilot boarded the ship at 0758 h. The vessel departed Wynyard Wharf with the last line away at 0821 h, beginning our sea passage. The rest of the day was spent in transit to Site U1527 (proposed Site NWC-1A).


Daily Science Report for 8 May 2018

Location: Freyberg Wharf B/Wynyard Wharf, Port of Auckland, New Zealand

Science Update: The Expedition 376 Curator gave a talk on curation, followed by an introduction to core sampling on the core deck. Half of the science party was given a core flow tour by the Expedition Project Manager (EPM). All laboratory groups convened for meetings with the IODP JRSO technical staff to learn their laboratories and start preparing associated instruments. The loading of supplies and hardware was completed, including the CDEX turbine-driven coring system. A news team of the New Zealand nationwide TV channel “Three” visited the vessel and was given a tour and interview by the Co-Chief Scientists and EPM. The ship shifted to Wynyard Wharf for fueling with the first line ashore at 1930 h.


Daily Science Report for 7 May 2018

Location: Freyberg Wharf B, Port of Auckland, New Zealand

Science Update: Today began with introductions of the Expedition 376 scientists and IODP JRSO technical staff, a talk on the expedition science objectives by the Co-Chief Scientists, and a presentation of the expedition work plan by the Expedition Project Manager. In the afternoon, after the Captain’s introduction and safety orientation, scientists presented their individual research goals. The day was concluded with an introduction to the descriptive information system, given by IODP JRSO staff to the core describers. Initial laboratory team meetings took place. The loading of supplies and essential hardware (e.g., reentry systems) continued all day. We are tentatively scheduled to depart at 0800 h on May 9.


Daily Science Report for 6 May 2018

Location: Freyberg Wharf B, Port of Auckland, New Zealand

Science Update: The Expedition 376 scientists boarded the vessel at 0900 h, got settled in their rooms, and were given a short introductory talk and presentation on information technology aboard the JOIDES Resolution. They started to connect their computers to the shipboard network and had their photos taken. At 1300 h, the science party convened for presentations about life at sea and general safety that were concluded by a ship safety tour. Seven PR tours of the vessel took place as scheduled. Loading and discharge of freight continued throughout the day.


Daily Science Report for 5 May 2018

Location: Freyberg Wharf B, Port of Auckland, New Zealand

Science Update: The Brothers Arc Flux Expedition 376 started at 0652 h with the first line ashore at the Freyberg Wharf B in Auckland. The Co-Chief Scientists, IODP JRSO staff, and two CDEX technical engineers moved onto the ship around 0900 h. IODP JRSO technical staff conducted crossover with their Expedition 375 counterparts. Expedition 375 and 376 Co-Chief Scientists’ interviews with Radio New Zealand and a tour of the ship took place as scheduled. Initial loading of incoming freight began.